OpenStudy (deadshot):

How do I find f(g(x)) if f(x) = 3x + 1 and g(x) = x^2 - 1?

OpenStudy (anonymous):

g(DeadShot) = DeadShot^2 - 1 isn't it?

OpenStudy (anonymous):

now what is g(f(x))?

OpenStudy (deadshot):

so f(g(x)) = 3(g(x)) + 1, right?

OpenStudy (anonymous):

YUP

OpenStudy (akashdeepdeb):

3(x) = 3x+1 So f(g(x)) = 3*g(x) + 1 [Now put g(x)'s value in it] Understood? :D

OpenStudy (deadshot):

\[f(g(x)) = 3*(x^2-1)+1\]

OpenStudy (amistre64):

f(x) tells us that the function f, is composed of x f(g) simply tells us that the function f, is composed of g, and we know how g is composed since they tell us that already

OpenStudy (deadshot):

So then \[f(g(x))= 3x^2-2\] or \[f(g(x))=3x^3-2\]?

OpenStudy (amistre64):

\[f(g(x)) = 3*(x^2-1)+1\] this does not produce an x^3 .... so

OpenStudy (deadshot):

so then the answer would be \[3x^2-2\]

OpenStudy (amistre64):

yes

OpenStudy (deadshot):

Thanks!

OpenStudy (amistre64):

youre welcome